# A stochastic volatility approximation for a tick-by-tick price model with mean-field interaction

### v2

@@ Abstract
We consider a tick-by-tick model of price formation, in which buy and sell orders are modeled as self-exciting point processes (Hawkes process), similar to the one in [Hoffmann, Bacry, Delattre, Muzy, Modelling microstructure noise with mutually exciting point processes, Quantitative Finance, 2013] and [El Euch, Fukasawa, Rosenbaum, The microstructural foundations of leverage effect and rough volatility, Finance and Stochastics, 2018]. We adopt an agent based approach by studying the aggregation of a large number of these point processes, mutually interacting in a mean-field sense.
The financial interpretation is that of an asset on which several labeled agents place buy and sell orders following these point processes, influencing the price. The mean-field interaction introduces positive correlations between order volumes coming from different agents that reflect features of real markets such as herd behavior and contagion. When the large scale limit of the aggregated asset price is computed, if parameters are set to a critical value, a singular phenomenon occurs: the aggregated model converges to a stochastic volatility model with leverage effect and faster-than-linear mean reversion of the volatility process.
@@ 1 Introduction
We follow here the general principle, inspired by Statistical Physics, of investigating to what extent macroscopic financial dynamics can be derived as limits of microscopic, possibly agent-based models [12, 16]. Moreover, inspired in particular by [20, 21, 36, 3, 38], the microscopic dynamics of buy and sell orders is modeled with Hawkes processes, which are self-exciting point processes. In particular, in the recent work [21], a rough version of the Heston model is derived as high-frequency limit of a market model in which price variations due to buy and sell orders are modeled as Hawkes processes with long memory. In addition to roughness, this provides a microscopic foundation of the leverage effect, i.e., the negative correlation between observed changes in stock returns and volatility. Unlike [21], where the microscopic model concerns the total flow of buy and sell orders, we introduce an agent based dynamics, by modeling the orders of each individual agent. The intensity according to which agents place buy and sell orders is a function of a time-weighted integral of the total past orders; this introduces a simple form of mean field interaction among agents. The simplicity lies in the fact that the total numbers of buy and sell orders are sufficient statistics: their aggregated dynamics can be expressed without referring to individual agents. This dimensional reduction appears in other classical mean field models, such as the mean field versions of the Voter Model, the Contact Process and the stochastic Ising model. As a consequence, the total numbers of buy and sell orders follows dynamical rules similar to those in [21], with the difference that we allow a nonlinear dependence of the intensities on the past orders. Besides this nonlinearity, that we discuss later, the agent based formulation offers at least two advantages. On one hand the hydrodynamic scaling needed for the limit is clearly linked to the size of agent’s population, providing a sound interpretation of the scaling regime to which the limit stochastic volatility model provides a good approximation to the microscopic dynamics. On the other hand, our model allows natural extensions, including inhomogeneities in the agent’s population (e.g. informed and uninformed agents) or metaorders modeling through high self-excitation of orders by the same agent; in the last case the aggregated numbers of buy and sell orders do not admit closed dynamical roles. These extensions will be discussed in Subsection 2.1.
@@
The nonlinearity in the dependence of the intensities on the past orders is the main feature of the model we propose. The linearity assumption is used e.g. in [21], and it is a key ingredient in the asymptotic analysis obtained there. We show that the introduction of nonlinearity, which could be motivated by saturation effects, significantly changes the asymptotic behavior. Relaxing linearity, moreover, forces us to use different arguments form those in [21]. We make a strict assumption on the memory function, that we suppose exponential; this allows a Markovian description of the pre-limit process and the use of tools from weak convergence to diffusion processes. Nonlinearity reveals the possibility of scaling properties which differ from the linear case, and naturally leads to quadratic mean reversion. This link between nonlinearity of the intensities in the pre-limit process and nonlinearity of the volatility mean reversion reveals further the effectiveness and flexibility of modeling with Hawkes processes.
@@
Finally, we mention again that our model keeps a relevant property of the linear models in [21], namely the leverage effect. Leverage effect refers to the widely documented negative correlation between observed changes in stock returns and volatility [22], see also [21, 39] and references therein. It is often incorporated in models for financial price dynamics through a negative correlation between the noises driving price and volatility, see [27, 32, 34, 48]. Leverage effect can be explained by the fact that when an asset price decreases, the ratio of the company’s debt with respect to the equity value becomes larger, and as a consequence volatility increases. Other “macroscopic” explanations from financial economics have been proposed [8, 14, 31, 25], as well as “psycological” explanations based on asymmetric responses by investors [30]. Here, as in [21], leverage effect observed on macroscopic quantities is a product of microstructural features, with in addition an explicit, stylized representation of interactions between agents.
@@ 2 Model description and main results
$m_{N}^{\pm}(t):=\frac{1}{N}\sum_{i=1}^{N}X^{\pm}_{i}(t).$
We will assume that the point processes $(N^{\pm}_{i})_{i=1}^{N}$ are conditionally independent given $(X^{\pm}_{i})_{i=1}^{N}$. Each “upward” jump process $N^{+}_{i}$ has a stochastic intensity given by
@@
where $a_{N}^{\pm},b_{N}^{\pm}$ are positive constants satisfying
@@
Note that $a_{N}^{\pm}$ is responsible for the value at time $t=0$ of the stochastic intensities $\lambda_{N}^{\pm}$, while $b_{N}^{\pm}$ represent stationary external signals. More irregular choices for $\nu_{N}$ are also possible.
@@
[Index and portfolio dynamics] A different possible interpretation of the aggregated price in (6) and its limit $(\pi_{t})_{t}$, is that of an index or portfolio composed of $i=1,\dots,N$ assets. The variations of the index are given by an average of the variations of the single components of the index, each represented by the sum of its positive and negative “micro” variations. The single assets have dynamics as in [21, 20], they are self exciting and are exciting each other, with jumps in one price exciting jumps in all the others through a mean-field interaction. The diffusive dynamics is seen at a slower time scale than the order book time scale where we see the jumps in the prices.
@@
[Multifractality - Multiscaling of moments] Let $(X_{t})_{t\geq 0}$ be a continuous-time martingale, having stationary increments; in financial applications this could be identified with the de-trended log-price of an asset, or the log-price with respect to the martingale measure used for pricing derivatives. We say that multiscaling of moments occurs if ${\mathbb{E}}\left(|X_{t+h}-X_{t}|^{q}\right)$ scales, in the limit as $h\downarrow 0$, as $h^{A(q)}$, with $A(q)=\frac{q}{2}$ for small $q$ but $A(q)<\frac{q}{2}$ beyond a certain threshold. The presence of multiscaling in financial time series is well established and various models have been proposed to capture it, most notably multi-fractal processes [10], see also [51, 44, 1]. In [17] we have shown that in a stochastic volatility model whose volatility $V_{t}$ solves the stochastic differential equation
@@ 2.1 Extensions of the model
In this section we illustrate two extensions of the model in Theorem 2.1. First we break the symmetry among agents, letting the model’s parameters to vary within the population. Later we introduce self-interaction: agent’s history has more impact on her own future actions than in that of other players, which we interpret as the fact that large trading orders are often split into smaller orders, executed over a certain time span, leading to an “apparent” high self-excitation [43]. This last model loses one of the key properties of the former two; namely that the aggregated means $m_{N}^{\pm}(t)$ evolve as a two dimensional Markov process. We state the theorems concerning the limiting behavior of these models. The proofs require rather mild modifications with respect to that of Theorem 2.1, and are given in Section 3.
@@
$\rho_{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{(\beta_{i},\gamma_{i})},$
@@
where $\rho$ is a probability on $1,+\infty)\times[0,1]$, with compact support, and $\alpha>0$. We also introduce the following notation: for $f:[1,+\infty)\times[0,1]\rightarrow{\mathbb{R}}$ set
$\overline{f(\beta,\gamma)}_{N}:=\frac{1}{N}\sum_{i=1}^{N}f(\beta_{i},\gamma_{i}),\hskip 28.45274pt\overline{f(\beta,\gamma)}:=\int f(\beta,\gamma)\rho(d\beta,d\gamma).$
@@
$\overline{f(\beta,\gamma)}=\lim_{N\rightarrow+\infty}\overline{f(\beta,\gamma)}_{N}$
if $f$ is continuous. We are now ready to state the extension of Theorem 2.1.
@@ 3.1 Preliminaries
are orthogonal martingales with conditional quadratic variations $\frac{1}{N}\int_{0}^{t}\lambda_{N}^{\pm}(s)ds$ (see [45] for the definition of conditional quadratic variation).
@@
which can be aggregated to obtain (18).
@@ 3.2 Definition and convergence of the volatility process
Then, we define $\tilde{Z}_{N}$ to be the process such that (22) holds for all $t\geq 0$ after replacing $Y_{N}$ and $Z_{N}$ by $\tilde{Y}_{N}$ and $\tilde{Z}_{N}$ (these processes appear as arguments of $\lambda^{\pm}$). This procedure is well defined since (22) can be seen as an equation for $Z_{N}$ given the process $Y_{N}$. Actually, for a given process $\tilde{Y}_{N}$ and after having extended by setting $f(x)=0$ for $x<0$, the corresponding solution $\tilde{Z}_{N}$ can be obtained pathwise, i.e. given the jumps of the processes $n_{i}^{\pm}$. Note that $\tilde{Y}_{N}(t)=Y_{N}(t)$ and $\tilde{Z}_{N}(t)=Z_{N}(t)$ up to the stopping time
@@
By Lemma 3.2 we have
where
$\tilde{\lambda}_{N}^{+}(t)=f\left(\frac{1}{\sqrt{N}}\tilde{Y}_{N}(t/\sqrt{N})+\frac{1}{\sqrt{N}}\frac{2\beta}{1+\beta}\tilde{Z}_{N}(t/\sqrt{N})\right)$
and
$\tilde{\lambda}_{N}^{-}(t)=f\left(\frac{1}{\sqrt{N}}\tilde{Y}_{N}(t/\sqrt{N})-\frac{1}{\sqrt{N}}\frac{2}{1+\beta}\tilde{Z}_{N}(t/\sqrt{N})\right)$
It follows that (see e.g. Theorem 31, Ch. II in [45])
@@
Observing that $\alpha-(1-\gamma)f^{\prime}(0)=(1+\beta\gamma)f^{\prime}(0)-(1-\gamma)f^{\prime}(0)=\beta(1+\gamma)f^{\prime}(0)>0$, the conclusion now follows from Lemma B.1.
@@
$\mathbb{E}\left[X_{N}^{d}(\varepsilon)\right]=\mathbb{E}\left[\tilde{Z}_{N}^{4}(\varepsilon)\right]\leq BN^{-\frac{1}{2}}\leq C_{1}N^{-\frac{1}{4}}=C_{1}\alpha_{N}^{-d}$
for some constants $B$ and $C_{1}$, so condition (78) is checked. For condition (79), we make estimates similar to those in Lemma 3.3. Recalling that
@@
We observe that for $t<\tau_{h}^{N}$ we have $\tilde{Y}_{N}(y)=Y_{N}(t)\leq h$ and, by (21), $|\tilde{Z}_{N}(t)|=|Z_{N}(t)|\leq C(\beta,\gamma)Y_{N}(t)\leq C(\beta,\gamma)h$. So condition (79) follows from (39), that $\alpha-f^{\prime}(0)(1-\gamma)>0$, and that $\sqrt{N}\geq n=N^{\frac{3}{16}}$. We now consider conditions (80) and (81). Note that, by (37),
$|f_{N}(s,u,\pm)|\leq\frac{C}{\sqrt{N}}\leq C_{4}\alpha_{N}^{-1}$
for some constants $C$ and $C_{4}$; moreover
$\begin{split}\int_{\mathcal{Y}}(f_{N}(t,y))^{2}A_{N}(t,dy)=&N^{\frac{3}{2}}\tilde{\lambda}^{+}(\sqrt{N}t)\left[\left(\tilde{Z}_{N}(s)+\frac{1-\gamma}{2\sqrt{N}}\right)^{2}-\tilde{Z}_{N}^{2}(s)\right]^{2}\\ &+N^{\frac{3}{2}}\tilde{\lambda}^{-}(\sqrt{N}t)\left[\left(\tilde{Z}_{N}(s)-\frac{1-\gamma}{2\sqrt{N}}\right)^{2}-\tilde{Z}_{N}^{2}(s)\right]^{2}\leq C_{5}\end{split}$
for some constant $C_{5}$, as, for some $C>0$,
@@
is a martingale with conditional quadratic variation $A_{N}$, i.e. $(M_{N}^{Y})^{2}-A_{N}$ is a martingale, as required in Theorem A.1.
@@
To prove (74) we show that all terms in the r.h.s of (50) converge to zero in probability after replacing $t$ by $t\wedge\tau^{N}_{h}$, with $\tau^{N}_{h}:=\inf\{t>0:Y_{N}(t)>h\}$. The first term in easy, since
@@
by (5). The estimate of the next two terms in (50) is based on the facts that, for $t\leq\tau^{N}_{h}$, we have that $\tilde{Z}_{N}(t)=Z_{N}(t)$, $Y_{N}(t)\leq h$ and, by (21)
$|Z_{N}(t)|\leq C(\beta,\gamma)Y_{N}(t)\leq C(\beta,\gamma)h.$
Thus
$\left|A(\beta,\gamma)f^{\prime\prime}(0)\int_{0}^{t\wedge\tau^{N}_{h}}Y_{N}(s)Z_{N}(s)ds+B(\beta,\gamma)f^{\prime\prime}(0)\int_{0}^{t\wedge\tau^{N}_{h}}Z_{N}^{2}(s)ds\right|\leq\mbox{constant}\cdot\int_{0}^{t}\tilde{Z}_{N}^{2}(s)\wedge C^{2}(\beta,\gamma)h^{2}ds$
which, by Lemma 3.3 and dominated convergence, converges to zero in $L^{1}$ and so in probability. For the remaining two terms in (50) we use again
$|Z_{N}(t)|\leq C(\beta,\gamma)Y_{N}(t),$
which implies that
$\int_{0}^{t\wedge\tau^{N}_{h}}R\left(\frac{1}{\sqrt{N}}Y_{N}\left(s\right)+\frac{1}{\sqrt{N}}\frac{2\beta}{1+\beta}Z_{N}\left(s\right)\right)\left(\frac{1}{\sqrt{N}}Y_{N}\left(s\right)+\frac{1}{\sqrt{N}}\frac{2\beta}{1+\beta}Z_{N}\left(s\right)\right)^{3}ds$
is bounded by a constant. This completes the proof on (74). The proof of (73) is similar but easier, as it requires only a first order expansion for $f$

### v3

@@ Abstract
We consider a tick-by-tick model of price formation, in which buy and sell orders are modeled as self-exciting point processes (Hawkes process), similar to the one in [Bacry, Delattre, Hoffmann, Muzy, Modelling microstructure noise with mutually exciting point processes, Quantitative Finance, 2013] and [El Euch, Fukasawa, Rosenbaum, The microstructural foundations of leverage effect and rough volatility, Finance and Stochastics, 2018]. We adopt an agent based approach by studying the aggregation of a large number of these point processes, mutually interacting in a mean-field sense.
The financial interpretation of the model is that of an asset on which several labeled agents place buy and sell orders following these point processes, influencing the price. The mean-field interaction introduces positive correlations between order volumes coming from different agents that reflect features of real markets such as herd behavior and contagion. When the large scale limit of the aggregated asset price is computed, if parameters are set to a critical value, a singular phenomenon occurs: the aggregated model converges to a stochastic volatility model with leverage effect and faster-than-linear mean reversion of the volatility process.
@@ 1 Introduction
We follow here the general principle, inspired by Statistical Physics, of investigating to what extent macroscopic financial dynamics can be derived as limits of microscopic, possibly agent-based models [14, 18]. Moreover, inspired in particular by [3, 22, 37, 4, 39], the microscopic dynamics of buy and sell orders is modeled with Hawkes processes, which are self-exciting point processes. In particular, in the recent work [22], a rough version of the Heston model is derived as high-frequency limit of a market model in which price variations due to buy and sell orders are modeled as Hawkes. In addition to roughness, this work provides a microscopic foundation of the leverage effect. Unlike [22], where the microscopic model concerns the total flow of buy and sell orders, we introduce an agent based dynamics, by modeling the orders of each individual agent. The intensity according to which agents place buy and sell orders is a function of a time-weighted integral of the total past orders; this introduces a simple form of mean field interaction among agents. The simplicity lies in the fact that the total numbers of buy and sell orders are sufficient statistics: their aggregated dynamics can be expressed without referring to individual agents. This dimensional reduction appears in other classical mean field models, such as the mean field versions of the Voter Model, the Contact Process and the stochastic Ising model. As a consequence, the total numbers of buy and sell orders follows dynamical rules similar to those in [22], with the difference that we allow a nonlinear dependence of the intensities on the past orders. Besides this nonlinearity, that we discuss later, the agent based formulation offers at least two advantages. On one hand the hydrodynamic scaling needed for the limit is clearly linked to the size of agent’s population, providing a sound interpretation of the scaling regime to which the limit stochastic volatility model provides a good approximation to the microscopic dynamics. On the other hand, our model allows natural extensions, including inhomogeneities in the agent’s population (e.g. informed and uninformed agents) or metaorders modeling through high self-excitation of orders by the same agent; in the last case the aggregated numbers of buy and sell orders do not admit closed dynamical roles. These extensions will be discussed in Subsection 2.1.
@@
The nonlinearity in the dependence of the intensities on the past orders is the main feature of the model we propose. The linearity assumption is used e.g. in [22], and it is a key ingredient in the asymptotic analysis obtained there. We show that the introduction of nonlinearity, which can be motivated by saturation effects, significantly changes the asymptotic behavior. The limit is still a stochastic volatility models, but with nonlinear mean reversion. Models with superlinear mean reversion have been proposed and supported by several authors [6, 47, 41, 7, 19]; to our knowledge this is the first research proposing a justification in the context of price dynamics. Relaxing linearity, moreover, forces us to use different arguments form those in [22]. We make a strict assumption on the memory function, that we suppose exponential; this allows a Markovian description of the pre-limit process and the use of tools from weak convergence to diffusion processes. Nonlinearity reveals the possibility of scaling properties which differ from the linear case, and naturally leads to quadratic mean reversion. This link between nonlinearity of the intensities in the pre-limit process and nonlinearity of the volatility mean reversion reveals further the effectiveness and flexibility of modeling with Hawkes processes.
@@
Finally, we mention again that our model keeps a relevant property of the linear models in [22], namely the leverage effect, which refers to the widely documented negative correlation between observed changes in stock returns and volatility [23], see also [22, 40] and references therein. It is often incorporated in models for financial price dynamics through a negative correlation between the noises driving price and volatility, see [28, 33, 35, 49]. Leverage effect can be explained by the fact that when an asset price decreases, the ratio of the company’s debt with respect to the equity value becomes larger, and as a consequence volatility increases. Other “macroscopic” explanations from financial economics have been proposed [9, 16, 32, 26], as well as “psycological” explanations based on asymmetric responses by investors [31]. Here, as in [22], leverage effect observed on macroscopic quantities is a product of microstructural features, with in addition an explicit, stylized representation of interactions between agents.
@@ 2 Model description and main results
$m_{N}^{\pm}(t):=\frac{1}{N}\sum_{i=1}^{N}X^{\pm}_{i}(t),$
that, as we will see in Lemma 3.1, form a two dimensional Markov process. We will assume that the point processes $(N^{\pm}_{i})_{i=1}^{N}$ are conditionally independent given $(X^{\pm}_{i})_{i=1}^{N}$. Each “upward” jump process $N^{+}_{i}$ has a stochastic intensity given by
@@
where $\delta_{0}$ denotes the Dirac measure and $a_{N}^{\pm},b_{N}^{\pm}$ are positive constants satisfying
@@
Note that $a_{N}^{\pm}$ is responsible for the value at time $t=0$ of the stochastic intensities $\lambda_{N}^{\pm}$, while $b_{N}^{\pm}$ represent stationary external signals.
@@
[Index and portfolio dynamics] A different possible interpretation of the aggregated price in (6) and its limit $(\pi_{t})_{t}$, is that of an index or portfolio composed of $i=1,\dots,N$ assets. The variations of the index are given by an average of the variations of the single components of the index, each represented by the sum of its positive and negative “micro” variations. Each individual asset has dynamics as in [22, 3]. The asset prices are self exciting and are exciting each other, with jumps in one price exciting jumps in all the others through a mean-field interaction. The diffusive dynamics is seen at a slower time scale than the order book time scale where we see the jumps in the prices.
@@
[Multifractality - Multiscaling of moments] Let $(X_{t})_{t\geq 0}$ be a continuous-time martingale, having stationary increments; in financial applications this could be identified with the de-trended log-price of an asset, or the log-price with respect to the martingale measure used for pricing derivatives. We say that multiscaling of moments occurs if ${\mathbb{E}}\left(|X_{t+h}-X_{t}|^{q}\right)$ scales, in the limit as $h\downarrow 0$, as $h^{A(q)}$, with $A(q)=\frac{q}{2}$ for small $q$ but $A(q)<\frac{q}{2}$ beyond a certain threshold. The presence of multiscaling in financial time series is well established and various models have been proposed to capture it, most notably multi-fractal processes [12], see also [52, 45, 1, 11]. In [19] we have shown that in a stochastic volatility model whose volatility $V_{t}$ solves the stochastic differential equation
@@ 2.1 Extensions of the model
In this section we illustrate two extensions of the model in Theorem 2.1. First we break the symmetry among agents, letting the model’s parameters to vary within the population. Later we introduce self-interaction: agent’s history has more impact on her own future actions than on that of other players, which we interpret as the fact that large trading orders are often split into smaller orders, executed over a certain time span, leading to an “apparent” high self-excitation [44]. This last model loses one of the key properties of the former two; namely that the aggregated means $m_{N}^{\pm}(t)$ evolve as a two dimensional Markov process (see Lemma 3.1). We state the theorems concerning the limiting behavior of these models. The proofs require rather mild modifications with respect to that of Theorem 2.1, and are given in Section 3.
@@
$\rho_{N}:=\frac{1}{N}\sum_{i=1}^{N}\delta_{(\beta^{N}_{i},\gamma^{N}_{i})},$
@@
where $\rho$ is a probability on $1,+\infty)\times[0,1]$, with compact support, and $\alpha>0$. We also introduce the following notation: for a function $g:[1,+\infty)\times[0,1]\rightarrow{\mathbb{R}}$ set
$\overline{g(\beta,\gamma)}_{N}:=\frac{1}{N}\sum_{i=1}^{N}g(\beta^{N}_{i},\gamma^{N}_{i}),\hskip 28.45274pt\overline{g(\beta,\gamma)}:=\int g(\beta,\gamma)\rho(d\beta,d\gamma).$
@@
$\overline{g(\beta,\gamma)}=\lim_{N\rightarrow+\infty}\overline{g(\beta,\gamma)}_{N}$
if $g$ is continuous. We are now ready to state the extension of Theorem 2.1.
@@ 3 Proofs
In this section we denote by $A,B,C$ constant independent of $N$, that may vary from line to line.
@@ 3.1 Preliminaries
$M_{N}^{\pm}(t):=\int_{[0,t]\times[0,+\infty)}{\bf 1}_{[0,\lambda_{N}^{\pm}(s-))}(u)\frac{1}{N}\sum_{i=1}^{N}\tilde{n}_{i}^{\pm}(ds,du)$
are orthogonal martingales with conditional quadratic variations $\frac{1}{N}\int_{0}^{t}\lambda_{N}^{\pm}(s)ds$ (see [46, Chapter III, Section 5, page 124] for the definition of conditional quadratic variation). Moreover the processes $m_{N}^{\pm}(t)$ form a two dimensional Markov process.
@@
which can be aggregated to obtain (18). Finally, the Markov property of $m_{N}^{\pm}(t)$ follows from the representation (18), the independence properties of Poisson Random Measures and the fact that the processes $\lambda_{N}^{\pm}(t)$ are deterministic functions of $m_{N}^{\pm}(t)$ (see (3) and (4)).
Remark 3.2.
Note that equations (20) show that the trajectories of $X^{\pm}_{i}$ are uniquely determined by the realization of the points of the Poisson Random Measures $n_{i}^{\pm}$. Indeed, $X_{i}^{\pm}(t)$ evolves deterministically until the moving intervals $0,\lambda_{N}^{\pm}(t))$ meet a point of one of the $n_{i}^{\pm}$, which causes a unit jump of $X_{i}^{\pm}(t)$.
@@ 3.2 Definition and convergence of the volatility process
Then, we define $\tilde{Z}_{N}$ to be the process such that (22) holds for all $t\geq 0$ after replacing $Y_{N}$ and $Z_{N}$ by $\tilde{Y}_{N}$ and $\tilde{Z}_{N}$ (these processes appear as arguments of $\lambda^{\pm}$). More precisely, given a realization of $n_{i}^{\pm}$, $i=1,\ldots,N$, which determines the trajectory of $Y_{N}$ and $\tilde{Y}_{N}$, define $\tilde{Z}_{N}$ as the process that satisfies the equation
where
$\tilde{\lambda}_{N}^{+}(t)=f\left(\frac{1}{\sqrt{N}}\tilde{Y}_{N}(t/\sqrt{N})+\frac{1}{\sqrt{N}}\frac{2\beta}{1+\beta}\tilde{Z}_{N}(t/\sqrt{N})\right)$ (24)
and
$\tilde{\lambda}_{N}^{-}(t)=f\left(\frac{1}{\sqrt{N}}\tilde{Y}_{N}(t/\sqrt{N})-\frac{1}{\sqrt{N}}\frac{2}{1+\beta}\tilde{Z}_{N}(t/\sqrt{N})\right)$ (25)
Similarly to what is done in Remark 3.2, the trajectory of $\tilde{Z}_{N}$ can be reconstructed pathwise, solving the deterministic part of the equation between jumps, and performing the jumps induced by the points of the $n_{i}^{\pm}$’s and the moving intervals $0,\tilde{\lambda}_{N}^{\pm}(s))$. Since the truncation may not guarantee the fact that the arguments of $f$ in (24) and (25) are positive, we extend $f$ by setting $f(x)=0$ for $x<0$. The relevant outcome of this truncation is the boundedness of $\tilde{Y}_{N}$, and the fact that $\tilde{Y}_{N}(t)=Y_{N}(t)$ and $\tilde{Z}_{N}(t)=Z_{N}(t)$ up to the stopping time
@@
It follows from (23) that (see e.g. Theorem 31, Ch. II in [46])
@@
From (7), observing that $\alpha-(1-\gamma)f^{\prime}(0)=(1+\beta\gamma)f^{\prime}(0)-(1-\gamma)f^{\prime}(0)=\beta(1+\gamma)f^{\prime}(0)>0$, the conclusion now follows from Lemma B.1.
@@
$\mathbb{E}\left[X_{N}^{d}(\varepsilon)\right]=\mathbb{E}\left[\tilde{Z}_{N}^{4}(\varepsilon)\right]\leq AN^{-\frac{1}{2}}\leq BN^{-\frac{1}{4}}=B\alpha_{N}^{-d}$
for some constants $A$ and $B$, so condition (82) is checked. For condition (83), we make estimates similar to those in Lemma 3.4. Recalling that
@@
We observe that for $t<\tau_{h}^{N}$ we have $\tilde{Y}_{N}(y)=Y_{N}(t)\leq h$ and, by (21),
$|\tilde{Z}_{N}(t)|=|Z_{N}(t)|\leq C(\beta,\gamma)Y_{N}(t)\leq C(\beta,\gamma)h.$ (42)
So condition (83) follows from (41), and the facts that $\alpha-f^{\prime}(0)(1-\gamma)>0$ and $\sqrt{N}\geq n=N^{\frac{3}{16}}$. We now consider conditions (84) and (85). Note that, by (39),
$|f_{N}(s,u,\pm)|\leq\frac{A}{\sqrt{N}}\leq B\alpha_{N}^{-1}$
for some constants $A$ and $B$; moreover
$\begin{split}\int_{\mathcal{Y}}(f_{N}(t,y))^{2}A_{N}(t,dy)=&N^{\frac{3}{2}}\tilde{\lambda}^{+}(\sqrt{N}t)\left[\left(\tilde{Z}_{N}(s)+\frac{1-\gamma}{2\sqrt{N}}\right)^{2}-\tilde{Z}_{N}^{2}(s)\right]^{2}\\ &+N^{\frac{3}{2}}\tilde{\lambda}^{-}(\sqrt{N}t)\left[\left(\tilde{Z}_{N}(s)-\frac{1-\gamma}{2\sqrt{N}}\right)^{2}-\tilde{Z}_{N}^{2}(s)\right]^{2}\leq A\end{split}$
for some constant $A$, as, for some $C>0$,
@@
and therefore is a martingale that, by (46), has conditional quadratic variation $A_{N}$, i.e. $(M_{N}^{Y})^{2}-A_{N}$ is a martingale, as required in Theorem A.1.
@@
To prove (78) we show that all terms in the r.h.s of (53) converge to zero in probability after replacing $t$ by $t\wedge\tau^{N}_{h}$, with $\tau^{N}_{h}:=\inf\{t>0:Y_{N}(t)>h\}$. The first term is easy, since
@@
by (5). The estimate of the next two terms in (53) is based on (42), which gives
$\left|A(\beta,\gamma)f^{\prime\prime}(0)\int_{0}^{t\wedge\tau^{N}_{h}}Y_{N}(s)Z_{N}(s)ds+B(\beta,\gamma)f^{\prime\prime}(0)\int_{0}^{t\wedge\tau^{N}_{h}}Z_{N}^{2}(s)ds\right|\leq\mbox{constant}\cdot\int_{0}^{t}\tilde{Z}_{N}^{2}(s)\wedge C^{2}(\beta,\gamma)h^{2}ds,$
where $C(\beta,\gamma)$ is here the constant in (42). This last term, by Lemma 3.4 and dominated convergence, converges to zero in $L^{1}$ and so in probability. For the remaining two terms in (53) we use again (21), which implies that
$\int_{0}^{t\wedge\tau^{N}_{h}}R\left(\frac{1}{\sqrt{N}}Y_{N}\left(s\right)+\frac{1}{\sqrt{N}}\frac{2\beta}{1+\beta}Z_{N}\left(s\right)\right)\left(\frac{1}{\sqrt{N}}Y_{N}\left(s\right)+\frac{1}{\sqrt{N}}\frac{2\beta}{1+\beta}Z_{N}\left(s\right)\right)^{3}ds\leq C$ (54)
for some constant $C$. This implies that the first cubic term in (53) goes to zero uniformly; the second term is similar and dealt with in the same way. This completes the proof of (78). The proof of (77) is similar but easier, as it requires only a first order expansion for $f$